OpenStudy (anonymous):

Solve each exponential equation. Give the exact value for x. (a) 5^x = 0.04 x = (b) 7^x = 10 x = (c) 8^x = 6.5 x = (d) 4^x = 5 x =

OpenStudy (anonymous):

I'll do one and then see if you can do the others: We have to take the log of both sides and then use our log laws to rearrange for x to be the subject. \[5^x=0.04\] \[log(5^x)=log(0.04)\] \[xlog(5)=log(0.04)\] \[x=\frac{log(0.04)}{log(5)}\] Does that make sense?

OpenStudy (anonymous):

It does make sense and I appreciate your assistance. :)

OpenStudy (anonymous):

You're welcome :P